# Towards the Worst-case Robustness of Large Language Models

### v1

@@ 6 Experiment
Table 3: The average certified length against adversarial suffix attack using Llama3-8B (Dubey et al., 2024).
@@
DiffTextPure (Absorb) 3.87 6.57 12.35 $\infty$
DiffTextPure (Uniform) 3.72 6.41 11.47 $\infty$
In this section, we demonstrate the experimental details of both empirical evaluation and certified robustness.
6.1 Empirical Evaluations
Settings. We conduct both black-box evaluations to demonstrate practical usage (Sec. E.3) and white-box evaluations (Table 1) to establish the upper bound of worst-case robustness. Following Zou et al. (2023); Jia et al. (2024); Liao & Sun (2024), we use the AdvBench dataset (Zou et al., 2023). We perform suffix attacks that append $d=20$ adversarial tokens as a suffix to the original request and optimize these appended tokens. We set $\beta=0.25$. Refer to Appendix E for other details.
Results. As demonstrated, all defenses achieve reasonable performance in black-box settings. This is because current optimization-based attacks, which do not utilize human-crafted prompts, transfer unreasonably poorly across different LLMs (Schaeffer et al., 2024), making black-box defenses much easier. However, these defenses all achieve nearly 0% white-box robustness. This is consistent with the previous experience in vision domains that only a few defenses that are not 0% robustness against white-box attacks (Athalye & Carlini, 2018; Athalye et al., 2018). As a randomization-based defense, DiffTextPure achieves non-trivial white-box robustness. This is partially because the stochasticity significantly interferes with white-box evaluations (Lee & Kim, 2023; Liu et al., 2024b) and partially because of the intrinsic robustness of such smoothing defenses (Carlini et al., 2023b; Cohen et al., 2019).
6.2 Certified Robustness
Settings. We still use the AdvBench dataset (Zou et al., 2023). Detectors are constructed by adjusting the prompt of the LLM (see Sec. E.1). This prompt is extremely conservative, ensuring that the detector achieves a 0% false positive rate on normal requests, like MT-Bench (Zheng et al., 2024a). Additionally, we perform certification on sentences containing more harmful keywords in Sec. E.7, using different purifiers in Sec. E.6, and ablation studies in Appendix E.
We focus on certifying safety against two types of attacks: In $\ell_{0}$ attack, we set $\beta=0.1$. We apply DiffTextPure to the entire sentence, thereby certifying the $\ell_{0}$ radius. In Suffix Attack, we set $\beta=0.25$. we first pad the input sentence with 50 arbitrary tokens and apply DiffTextPure to all except the first $k$ tokens. We purify the diffused tokens conditioned on the undiffused tokens. If the suffix part is not fully included in DiffTextPure, we return a certified robustness of 0. Otherwise, we certify the $\ell_{0}$ radius excluding the first $k$ tokens, thereby covering the suffix part. We also discuss additional settings in Sec. F.4.
Results. As shown in Table 2, for $\ell_{0}$ attacks, we achieve a certified radius of $2.02$. The better the base model, the higher the true positive rate, and thus, the higher the certified radius. For the AdvBench dataset, the theoretical upper limit—Bayesian bound—of the certified method is close to the real limit, the trivial bound, demonstrating that the certified bound closely approximates real robustness. However, this does not always hold true for datasets containing longer requests (Sec. E.7). This may require a fundamental improvement on the randomized smoothing paradigm, e.g., relying on more neural network-dependent variables rather than a single $p_{A}$. For adversarial suffix attacks, we achieve an average certified radius of $6.41$ (with $\beta=0.25$), while practical settings focus on suffix lengths of $20$. This demonstrates that it is relatively easy to obtain a certified radius with strong practical significance in the suffix attack settings due to its simplicity (Kumar et al., 2023).
Smoothness-utility trade-off. As the probability of perturbing a token $\beta$ approaches 1, the distribution of the diffused samples becomes identical for both benign and adversarial inputs. In this scenario, the base model is unable to distinguish whether the noisy examples come from benign or adversarial inputs. As a result, the function $g$ becomes overly smooth, returning a constant output regardless of the input. Since we require a false positive rate of 0, in the $\ell_{0}$ attack setting, this directly leads to a certified radius of 0. In the suffix attack setting, the detector relies entirely on the first $k$ non-diffused tokens, causing the certified radius to be either 0 or $\infty$. A certified radius of $\infty$ is undesirable because adding several tokens can indeed alter the semantics of an input sentence (Sec. F.3). Typically, we choose $\beta=0.25$, as this value avoids masking key information in the sentence and prevents oversmoothing.
@@ 7 Conclusion
In this work, we investigate the worst-case robustness of large language models. We upper bound the worst-case robustness of previous defenses by proposing a strong adaptive attack that strictly ensures the consistency in tokenization between optimization and inference. We also lower bound the worst-case robustness of all randomization-based defenses by reducing the functional optimization to a fractal knapsack problem or 0-1 knapsack problem. We reduce the time complexity of certification exponentially for specific kernels, such as uniform kernels or absorbing kernels, by categorizing each item based on its value-to-weight ratio and pre-computing the number of items in each category. We also provide theoretical analysis on the relationship between certified robustness, smoothing distribution, and vocabulary size, and upper bound the certified lower bound by Bayesian error, offering insights into the upper limits of certified methods. See Appendix H for key takeaways.
@@ B.1 More Related Work on Jailbreak Attacks and Defenses
The jailbreak attack on LLMs primarily refers to inducing LLMs into generating harmful content that is unsafe or toxic to society (Chao et al., 2024; Zhou et al., 2024b). To achieve this goal, malicious attackers can craft jailbreaking prompts through manual design, optimization, or train a generative model. Manual-designed jailbreak prompts leverage heuristic perspectives like data distribution (Wei et al., 2023b; Deng et al., 2023; Wei et al., 2023a), psychology insights (Shen et al., 2024b; Zeng et al., 2024; Shen et al., 2024a; Li et al., 2023b) or cipher encoding (Yuan et al., 2023; Handa et al., 2024) to achieve this goal. Optimization-based attacks extend from manually designing by optimizing an adversarial prompt with certain loss functions, where they can optimize a prefix or suffix (Zou et al., 2023; Liu et al., 2023; Jia et al., 2024; Zhang & Wei, 2025; Li et al., 2024a), or directly refine the jailbreaking prompt (Dong et al., 2023; Chen et al., 2024c; Zheng et al., 2024b; Chao et al., 2023). Besides, a thread of work toward fitting the jailbreak prompt distribution with a generative model (Liao & Sun, 2024; Kumar et al., 2024; Paulus et al., 2024; Basani & Zhang, 2024), effectively increasing the attack efficiency. Notably, there are also fine-tuning-based attacks that directly manipulate the alignment instead of designing prompts (Qi et al., 2023; Yang et al., 2023; Zhang et al., 2024b), posing another safety threat to LLMs.
From the defense perspective, various methods are proposed at different stages of generation. Pre-processing defenses are designed to detect potential jailbreaking prompts, typically aimed at adversarial suffix-based attacks that cause significantly high perplexity (Jain et al., 2023; Alon & Kamfonas, 2024). Besides, prompt-based defenses add safety tokens during generation, which are manually designed (Wei et al., 2023b; Xie et al., 2023) or optimized (Mo et al., 2024; Zhou et al., 2024a). Finally, post-processing defenses detect jailbreaking with hidden spaces (Li et al., 2025; Galinkin & Sablotny, 2024) or toxicity detection (Wang et al., 2023c; Hu et al., 2024; Wang et al., 2024).
@@ B.2 Adversarial Attacks and Defenses on Text Domain
Textual adversarial attacks (Morris et al., 2020; Wang et al., 2019b; Han et al., 2022) extend adversarial examples from vision space to discrete text space. Thus, a major challenge of textual attacks is the optimization process on discrete tokens, which include character, word, or sentence-level attacks. For instance, word-level attacks replace critical tokens with semantically similar alternatives to evade detection (Jin et al., 2020; Zang et al., 2019), while character-level attacks insert misspellings or Unicode artifacts to bypass filters (Ebrahimi et al., 2018; Rocamora et al., 2024). Recent advances also employ generative models to automate the creation of adversarial examples (Ren et al., 2020; Li et al., 2023a), producing fluent but malicious inputs that align with natural language patterns. These attacks highlight the vulnerability of text-based systems to carefully crafted inputs, even when perturbations are imperceptible to humans.
Defending against textual adversarial attacks also requires addressing the discrete nature of language. Adversarial training (Xiao et al., 2018), which incorporates perturbed examples during model optimization, remains a cornerstone for improving the robustness of language models (Wang et al., 2019a; Gao et al., 2023). A series of certified defenses with randomized smoothing techniques provide probabilistic guarantees against textual bounded perturbations (Jia et al., 2019; Wang et al., 2021) was also proposed. The evolving landscape of text-domain adversarial robustness underscores the need for defenses that generalize across attack vectors while preserving linguistic integrity. However, these defenses and certifications are limited to conventional language models like sentence classifiers, yet the certified robustness of large generative models remains unexplored.
@@ B.3 Diffusion Models for Adversarial Robustness
Diffusion models (Song et al., 2021; Dhariwal & Nichol, 2021) have achieved notable success in defending against visual adversarial examples (Nie et al., 2022; Wang et al., 2022; Li et al., 2024b; Xiao et al., 2023; Zhang et al., 2023; Carlini et al., 2023b). In particular, they are widely used as a plug-and-play purification method, named DiffPure, making them suitable for commercial models (Zhang et al., 2024a). As illustrated in Figure 1, given a model to be protected model, $f$, and a diffusion denoiser $D$, DiffPure involves two main steps: First, it adds Gaussian noise with variance $\sigma_{\tau}^{2}$ to the input images, and then denoising these noisy images using the diffusion model $D$.
Intuitively, the norm of the added Gaussian noise is much larger than that of the adversarial perturbations, effectively washing out the adversarial nature of the small-norm perturbations (Nie et al., 2022). Theoretically, this procedure not only increases the log-likelihood of input images, pushing them back from out-of-distribution to in-distribution (Nie et al., 2022; Xiao et al., 2023), but also implicitly constructs a smooth classifier $g(\bm{x})=\mathbb{E}_{\bm{x}_{\tau}\sim\mathcal{N}(\bm{x},\sigma_{\tau}^{2}\bm{I})}[f(D(\bm{x}_{t}))]$. The mathematical properties of this classifier have been extensively studied, providing theoretical proof on whether adversarial examples can exist within certain neighborhoods (Carlini et al., 2023b; Xiao et al., 2023; Chen et al., 2024b; Zhang et al., 2023).
@@ B.4 More Related Work on Certified Robustness
Certified robustness by masking. Certified robustness through masking has been extensively studied in previous work (Zeng et al., 2023; Levine & Feizi, 2020; Moon et al., 2023; Zhang et al., 2019) in both text and image domains (e.g., partitioning images into patches and masking them). The certification approach for DiffTextPure-Absorb differs slightly from these works, as tokens are masked with a probability rather than at a fixed ratio, leading to a much more neat result, as shown in Sec. 5.3. Zeng et al. (2023) suggest that this certified lower bound can be improved by introducing an auxiliary variable. However, their approach does not incorporate hypothesis testing or account for type-one error in estimating this auxiliary variable. For randomized smoothing via masking, it is obvious that this bound is tight as there exists a worst-case $f$ that fails entirely on region $L_{1}$. When fixing their bound with hypothesis testing using Bonferroni correction, it is clear that this produces the same result.
@@
Certified robustness by random perturbing words. Jia et al. (2019) uses interval bounds propagation to propagate the activation bounds to the final layers. These methods currently are not scalable to large models. On the contrary, we adopt randomized smoothing, a model-agnostic certification approach, which is thus more scalable.
@@
Universal certification. Lee et al. (2019) also establish a lower bound when smoothing a pre-trained model with randomly perturbed words, but there are several key differences compared to our work. First, we demonstrate that the certified robustness problem can be formulated as a Fractal Knapsack problem, making the approach more intuitive and easier. Second, we show that this can be further improved when the base model $f$ is a hard function, which becomes a 0-1 Knapsack problem and can obtain a stronger result using dynamic programming. What’s more, we greatly simplify the problem by showing that only the different part needs to be considered (see Sec. 5.4), which significantly streamlines the computation of the value-to-weight ratio (see Theorem 5.6). Finally, we show that the uniform kernel reduces to the absorbing kernel when $|\mathcal{V}|\to\infty$, i.e., Figure 2(c) gradually becomes Figure 2(a), giving more theoretical insights.
@@
Certified robustness using synonyms substitution. Ye et al. (2020) perturbs words into synonyms (including the original word) with the same probability to achieve certified robustness against word substitution attacks using synonyms. This certified bound closely resembles our DiffTextPure-Absorb method. Specifically, for any perturbed sentence $\bm{z}$, either it cannot result from perturbing the natural or adversarial sentence (trading rate of 0), or it is derived from both with the same probability (trading rate of 1). Consequently, the procedure of certifying using this synonym distribution is the same as that of our absorbing kernel. This approach cannot be generalized to certify word substitution attacks beyond synonyms, as perturbing uniformly into each word in the whole vocabulary with the same probability would completely disrupt the semantics.
@@
Certified robustness for large language models. Kumar et al. (2023) first certify large language models against suffix attacks and insertion attacks by randomly deleting tokens. In our notation, they set $p(\bm{z}|\bm{x})$ as a uniform distribution over sentences that have deleted fewer than $k$ tokens from $\bm{x}$, and they set the threshold to infinitesimally small, i.e., as long as there is one harmful $\bm{z}$, they classify $\bm{x}$ as harmful. Therefore, their certified accuracy is exactly the empirical accuracy of detectors on the original text. However, this heuristic method has a very large false positive rate, because a benign sentence can also be deleted into a harmful one. For example, “How to prevent making a bomb” can be deleted into “How to make a bomb.” From our framework, we can easily propose a fix for their method by setting the threshold $\tau$ to a value greater than zero.
Ji et al. (2024) also use a language model as a purifier and certify large language models on classification tasks, but not on safety tasks. The most significant difference between safety and classification is that safety requires a very low false positive rate, while classification does not. Certifying safety, therefore, is much more difficult than certifying classification.
@@
Robey et al. (2023) propose smoothing a language model by randomly perturbing each character, rather than tokens. They also do not certify their defense. Their theorem is based on an assumption they define themselves, called k-stable, which states that perturbing $k+1$ characters would result in a change. This assumption is impractical and implicitly implies robustness. In this work, we do not make any such assumptions. Instead, we certify each input $\bm{x}$ independently, rather than relying on a distribution.

### v2

@@ 6 Experiment
In this section, we demonstrate the experimental details of both empirical evaluation for upper bounding worst-case robustness and theoretical certification for lower bounding worst-case robustness.
6.1 Empirical Evaluations
Settings. We conduct both black-box evaluations to demonstrate practical usage (Sec. G.3) and white-box evaluations (Table 1) to establish the upper bound of worst-case robustness. Following [10, 17, 57], we use the AdvBench dataset [10]. We perform suffix attacks that append $d=20$ adversarial tokens as a suffix to the original request and optimize these appended tokens. We set $\beta=0.25$. Refer to Appendix G for other details.
Results. As shown in Sec. G.3, all defenses achieve reasonable performance in black-box settings, demonstrating their high practicality. For white-box settings, see Sec. 3 for details.
6.2 Certified Robustness
Settings. We use the AdvBench dataset [10] to evaluate certified lower bounds for three previous empirical defenses: uniform kernel [30, 20], absorbing kernel [56, 55, 29], and SmoothLLM [31] (i.e., a uniform kernel applied to each character instead of each token). Note that the results for SmoothLLM presented in this paper certify character-level robustness rather than token-level robustness. We welcome contributions of new stochastic defenses for large language models and will provide their lower bounds using our solver, incorporating results into the main table.
We focus on certifying safety against two types of attacks. In the $\ell_{0}$ attack, we set $\beta=0.1$ and apply these defenses to the entire sentence, thereby certifying the $\ell_{0}$ radius. In the Suffix Attack, we set $\beta=0.25$, pad the input sentence with 50 arbitrary tokens, and apply these defenses to all tokens except the first $k$ tokens. Safety detectors are constructed by adjusting the prompt of the LLM (see Sec. G.1). This prompt is highly conservative, ensuring a 0% false positive rate on normal requests across datasets, including [58, 59, 60, 61].
Baseline. For $\ell_{0}$ attacks, certified radii cannot be arbitrarily large. For example, "how to make a bomb" can become "how to make a cake" by changing one token, thus the certified radius of this sentence cannot exceed 0. The “Human” baseline serves as an upper bound for the certified radius; see Sec. I.2 for details. For suffix attacks, we compare randomized smoothing with the method of Kumar et al. [62], which deletes the suffix and evaluates the detector on the resulting sentence. Consequently, the certified robustness equals the clean accuracy (i.e., 1), and the certified radius is infinite. All randomized smoothing methods degrade to Kumar et al. [62] when $\beta\to 1$.
Results. As shown in Table 2, for $\ell_{0}$ attacks, we achieve a certified radius of $2.02$. The better the base model, the higher the true positive rate, and thus, the higher the certified radius. For the AdvBench dataset, the obtained theoretical lower bound is close to the human performance. However, this does not always hold true, especially for datasets containing longer requests (Sec. G.7). This may require a fundamental improvement on the randomized smoothing paradigm, e.g., relying on more neural network-dependent variables rather than a single $p_{A}$. For adversarial suffix attacks, we achieve an average certified radius of $6.41$ (with $\beta=0.25$), while practical settings focus on suffix lengths of $20$. This demonstrates that it is relatively easy to obtain a certified radius with strong practical significance in the suffix attack settings due to its simplicity [62].
Table 3: The average certified length against suffix attack using Llama-3-8B [3].
@@ 6.2 Certified Robustness
Absorb 3.87 6.57 12.35 $\infty$
Uniform 3.72 6.41 11.47 $\infty$
SmoothLLM 2.93 5.26 7.13 $\infty$
Kumar et al. [62] $\infty$ $\infty$ $\infty$ $\infty$
Smoothness-utility Trade-off. As $\beta$ approaches 1, the distribution of diffused samples becomes identical for both benign and adversarial inputs. In this scenario, the base model cannot distinguish whether noisy examples originate from benign or adversarial inputs. Consequently, the function $g$ becomes overly smooth, producing a constant output regardless of the input. Since we require a false positive rate of 0, in the $\ell_{0}$ attack setting, this directly results in a certified radius of 0. In the suffix attack setting, the detector relies solely on the prefix, leading to a certified radius of either 0 or $\infty$, and all randomized smoothing methods degrade to Kumar et al. [62]. A certified radius of $\infty$ may be undesirable, as adding a few tokens can significantly alter the semantics of an input sentence (see Sec. H.3). Typically, we choose $\beta=0.25$, as this value avoids masking critical information in the sentence and prevents oversmoothing.
@@ 7 Conclusion
In this work, we investigate the worst-case robustness of large language models. We upper bound the worst-case robustness of previous defenses by proposing a strong adaptive attack that strictly ensures the consistency in tokenization between optimization and inference. We also lower bound the worst-case robustness of all randomization-based defenses by reducing the functional optimization to a fractal knapsack problem or 0-1 knapsack problem. We conduct two case studies on smoothing the distribution of the diffusion models and masked generation, analyze their certified lower bound and clean accuracy, demonstrating their relationship. We also provide theoretical analysis on the relationship between certified robustness, smoothing distribution, and vocabulary size, and upper bound the certified lower bound by Bayesian error, offering insights into the upper limits of certified methods. See Appendix K for key takeaways.
@@ B.1 More Related Work on Jailbreak Attacks and Defenses
The jailbreak attack on LLMs primarily refers to inducing LLMs into generating harmful content that is unsafe or toxic to society [63, 64]. To achieve this goal, malicious attackers can craft jailbreaking prompts through manual design, optimization, or train a generative model. Manual-designed jailbreak prompts leverage heuristic perspectives like data distribution [36, 65, 12], psychology insights [66, 67, 68, 69] or cipher encoding [70, 71] to achieve this goal. Optimization-based attacks extend from manually designing by optimizing an adversarial prompt with certain loss functions, where they can optimize a prefix or suffix [10, 39, 17, 72, 73], or directly refine the jailbreaking prompt [74, 75, 76, 11, 77]. Besides, a thread of work toward fitting the jailbreak prompt distribution with a generative model [57, 78, 79, 80], effectively increasing the attack efficiency. Notably, there are also fine-tuning-based attacks that directly manipulate the alignment instead of designing prompts [81, 82, 83], posing another safety threat to LLMs.
From the defense perspective, various methods are proposed at different stages of generation. Pre-processing defenses are designed to detect potential jailbreaking prompts, typically aimed at adversarial suffix-based attacks that cause significantly high perplexity [48, 84]. Besides, prompt-based defenses add safety tokens during generation, which are manually designed [36, 85] or optimized [41, 86]. Finally, post-processing defenses detect jailbreaking with hidden spaces [87, 88] or toxicity detection [89, 90, 91].
@@ B.2 Adversarial Attacks and Defenses on Text Domain
Textual adversarial attacks [92, 93, 94] extend adversarial examples from vision space to discrete text space. Thus, a major challenge of textual attacks is the optimization process on discrete tokens, which include character, word, or sentence-level attacks. For instance, word-level attacks replace critical tokens with semantically similar alternatives to evade detection [55, 95], while character-level attacks insert misspellings or Unicode artifacts to bypass filters [96, 97]. Recent advances also employ generative models to automate the creation of adversarial examples [98, 99], producing fluent but malicious inputs that align with natural language patterns. These attacks highlight the vulnerability of text-based systems to carefully crafted inputs, even when perturbations are imperceptible to humans.
Defending against textual adversarial attacks also requires addressing the discrete nature of language. Adversarial training [100], which incorporates perturbed examples during model optimization, remains a cornerstone for improving the robustness of language models [101, 102]. A series of certified defenses with randomized smoothing techniques provide probabilistic guarantees against textual bounded perturbations [30, 103] was also proposed. The evolving landscape of text-domain adversarial robustness underscores the need for defenses that generalize across attack vectors while preserving linguistic integrity. However, these defenses and certifications are limited to conventional language models like sentence classifiers, yet the certified robustness of large generative models remains unexplored.
@@ B.3 Diffusion Models for Adversarial Robustness
Diffusion models [104, 105] have achieved notable success in defending against visual adversarial examples [106, 107, 108, 109, 110, 111]. In particular, they are widely used as a plug-and-play purification method, named DiffPure, making them suitable for commercial models [112]. As illustrated in Figure 2, given a model to be protected model, $f$, and a diffusion denoiser $D$, DiffPure involves two main steps: First, it adds Gaussian noise with variance $\sigma_{\tau}^{2}$ to the input images, and then denoising these noisy images using the diffusion model $D$.
Intuitively, the norm of the added Gaussian noise is much larger than that of the adversarial perturbations, effectively washing out the adversarial nature of the small-norm perturbations [106]. Theoretically, this procedure not only increases the log-likelihood of input images, pushing them back from out-of-distribution to in-distribution [106, 109], but also implicitly constructs a smooth classifier $g(\bm{x})=\mathbb{E}_{\bm{x}_{\tau}\sim\mathcal{N}(\bm{x},\sigma_{\tau}^{2}\bm{I})}[f(D(\bm{x}_{t}))]$. The mathematical properties of this classifier have been extensively studied, providing theoretical proof on whether adversarial examples can exist within certain neighborhoods [111, 109, 113, 110].
@@ B.4 More Related Work on Certified Robustness
Certified robustness by masking. Certified robustness through masking has been extensively studied in previous work [29, 114, 115, 116] in both text and image domains (e.g., partitioning images into patches and masking them). The certification approach for DiffTextPure-Absorb differs slightly from these works, as tokens are masked with a probability rather than at a fixed ratio, leading to a much more neat result, as shown in Sec. 5.1. Zeng et al. [29] suggest that this certified lower bound can be improved by introducing an auxiliary variable. However, their approach does not incorporate hypothesis testing or account for type-one error in estimating this auxiliary variable. For randomized smoothing via masking, it is obvious that this bound is tight as there exists a worst-case $f$ that fails entirely on region $L_{1}$. When fixing their bound with hypothesis testing using Bonferroni correction, it is clear that this produces the same result.
@@
Certified robustness by random perturbing words. Jia et al. [30] uses interval bounds propagation to propagate the activation bounds to the final layers. These methods currently are not scalable to large models. On the contrary, we adopt randomized smoothing, a model-agnostic certification approach, which is thus more scalable.
@@
Universal certification. Lee et al. [117] also establish a lower bound when smoothing a pre-trained model with randomly perturbed words, but there are several key differences compared to our work. First, we demonstrate that the certified robustness problem can be formulated as a Fractional Knapsack problem, making the approach more intuitive and easier. Second, we show that this can be further improved when the base model $f$ is a hard function, which becomes a 0-1 Knapsack problem and can obtain a stronger result using dynamic programming. What’s more, we greatly simplify the problem by showing that only the different part needs to be considered (see Sec. 5.2), which significantly streamlines the computation of the value-to-weight ratio (see Theorem 5.4). Finally, we show that the uniform kernel reduces to the absorbing kernel when $|\mathcal{V}|\to\infty$, i.e., Figure 1(c) gradually becomes Figure 1(a), giving more theoretical insights.
@@
Certified robustness using synonyms substitution. Ye et al. [28] perturbs words into synonyms (including the original word) with the same probability to achieve certified robustness against word substitution attacks using synonyms. This certified bound closely resembles our DiffTextPure-Absorb method. Specifically, for any perturbed sentence $\bm{z}$, either it cannot result from perturbing the natural or adversarial sentence (trading rate of 0), or it is derived from both with the same probability (trading rate of 1). Consequently, the procedure of certifying using this synonym distribution is the same as that of our absorbing kernel. This approach cannot be generalized to certify word substitution attacks beyond synonyms, as perturbing uniformly into each word in the whole vocabulary with the same probability would completely disrupt the semantics.
@@
Certified robustness for large language models. Kumar et al. [62] first certify large language models against suffix attacks and insertion attacks by randomly deleting tokens. In our notation, they set $p(\bm{z}|\bm{x})$ as a uniform distribution over sentences that have deleted fewer than $k$ tokens from $\bm{x}$, and they set the threshold to infinitesimally small, i.e., as long as there is one harmful $\bm{z}$, they classify $\bm{x}$ as harmful. Therefore, their certified accuracy is exactly the empirical accuracy of detectors on the original text. Since it is extremely easy to achieve 100% TPR on clean data, one will definitely get 100% certified accuracy and +$\infty$+ ∞ certified radius using Kumar et al. [62]. All the randomized smoothing methods degrade to Kumar et al. [62] against suffix attacks when $\beta\to 1$.
@@
Robey et al. [31] propose smoothing a language model by randomly perturbing each character, rather than tokens. They also do not certify their defense. Their theorem is based on an assumption they define themselves, called k-stable, which states that perturbing $k+1$ characters would result in a change. This assumption indeed already implicitly implies robustness. In this work, we do not make any such assumptions. Instead, we certify each input $\bm{x}$ independently, rather than relying on a distribution.